The table above summarizes the results of a replica-plating experiment where a mixture of 347 E. coli strains of different genotypesMedium Number of colonies that grew 347 Complete medium (contains all amino acids, vitamins, and a variety of carbon sources) were tested for their ability to growth on a variety of media. Based on the data presented in the table above: A. how many colonies are met(+) arg(-)? B. how many colonies are amp(R) met(-) arg(+)? C. how many colonies are amp(R) met(-) arg(-)? D. how many colonies are amp(S) met(+) arg(+)? can you give a detail explanation? I keep finding different answers and I don't find the logic.
Answer #1

Answer:

\rightarrowA. 301-261=40

\rightarrowB.57-51=6

\rightarrowC.4

\rightarrowD.261-51=210

                                                                     Please Rate My Answer......Thank......u...